# How Do Recommendation Models Amplify Popularity Bias? An Analysis from the Spectral Perspective

### v3

@@ 3.2 Popularity Bias Amplification Effect
As the embedding dimension $d$ decreases, the relative prominence of the principal singular value increases ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}\sigma_{k}^{2}\uparrow$) and the recommendation increasingly favors popular items.
@@
3.2.2 Theoretical Analyses. In this subsection, we focus on establishing a theoretical relationship between singular values and the ratio of popular items in recommendations. For readers interested in the theoretical support of the impact of gradient optimization, we refer them to the Appendix A.4, which are relatively straightforward by invoking recent gradient theory [15, 48]. For convenience, our analysis here concentrates on the ratio of the most popular item in top-1 recommendations. We have:
@@
$\small\eta\geq\frac{1}{n}\phi\left({\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}(\frac{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}{{{\sigma_{1}}}}-1)}\right)$ (3)
@@
where $\phi(a)=\sum_{u\in\mathcal{U}}{\mathbf{I}[p_{1u}>a]}$ is an inverse cumulative function calculating the number of elements $p_{1u}$ in the left principal singular vector $\bf p_{1}$ exceeding a given value $a$, and the function $\mathbf{I}[.]$ signifies an indicator function.
@@
The detailed proof is available in Appendix A.3. This theorem vividly showcases the influence of dimension collapse on popularity bias. Essentially, as dimension collapse intensifies the relative prominence of the principle singular value (i.e., $\frac{{{\sigma_{1}}}}{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}\uparrow$), the input of the function $\phi(.)$ decreases (i.e., $\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}(\frac{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}{{{\sigma_{1}}}}-1)\downarrow$). Given the monotonically decreasing nature of $\phi(.)$, dimension collapse thus escalates the ratio of most popular items in recommendations. Interestingly, the theorem illustrates the impact of a long-tailed distribution on popularity bias. A larger $\alpha$ (indicating a more skewed item popularity distribution) decreases the value of $\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}$, further elevating the lower bound of the ratio, intensifying bias.
4 Proposed Method
4.1 ReSN: Regulation with Spectral Norm
@@
The above analyses elucidate the essence of the popularity bias amplification — the undue influence of the principal spectrum. To counteract this, we propose ReSN with leveraging Spectral Norm Regularizer to penalize the magnitude of principal singular value:
${\mathcal{L}_{{\mathop{\rm Re}\nolimits}SN}}={\mathcal{L}_{R}}({\bf{Y}},{\bf{\hat{Y}}})+\beta||{\bf{\hat{Y}}}||_{2}^{2}$ (4)
@@
where ${\mathcal{L}_{R}}({\bf{Y}},{\bf{\hat{Y}}})$ is original recommendation loss, and ||.$||.||_{2}$ denote the spectral norm of a matrix measuring its principle singular value; $\beta$ controls the contribution from the regularizer.
@@
However, there are practical challenges: 1) the $n\times m$ dimensional matrix $\bf{\hat{Y}}$ can become exceptionally large, often comprising billions of entries, making direct calculations computationally untenable; 2) Existing methods to determine the gradient of the spectral norm are iterative [60, 52], which further adds computational overhead.
@@
To circumvent these challenges, we make two refinements: Firstly, given the alignment of the principal singular vector $\mathbf{q}_{1}$ with item popularity $\mathbf{r}$, the calculating of the spectral norm can be simplified into: $||{\bf{\hat{Y}}}||_{2}^{2}=||{\bf{\hat{Y}}}{{\bf{q}}_{1}}||^{2}\approx||{\bf{\hat{Y}r}}||^{2}/||\mathbf{r}||^{2}$, where ||.||||.||| | . | | denotes the L2-norm of a vector. It transforms the calculation of the complex spectral norm of a matrix to a simple L2-norm of a vector, avoiding iterative algorithms by leveraging the singular vector property. Further, the item popularity $\bf r$ can be quickly computed via ${\bf{r}}={\mathbf{Y}^{\top}}{{\bf{e}}}$, where ${\bf{e}}$ represents a $n$-dimension vector filled with ones. Secondly, we exploit the low-rank nature of the matrix $\bf{\hat{Y}}$. For models based on embeddings, $\bf{\hat{Y}}$ can be expressed as ${\bf{\hat{Y}}}=\mu({\bf{U}}{{\bf{V}}^{\top}})$, where ${\bf{U}}$ and ${{\bf{V}}}$ represent the embeddings associated with users and items, respectively, and $\mu(.)$ designates an activation function. Our approach turns to penalize the spectral norm of the matrix before the introduction of the activation function. This is motivated by the ease of computation: $||{\bf{U}}{{\bf{V}}^{\top}}||_{2}^{2}=||{\bf{U}}({{\bf{V}}^{\top}}{{{\bf{\tilde{q}}}}_{1}})||^{2}$, where ${{{\bf{\tilde{q}}}}_{1}}$ denotes the right principal vector of the matrix ${\bf{U}}{{\bf{V}}^{\top}}$. By adopting this method, we circumvent the computationally-intensive task of processing the entire matrix $\bf{\hat{Y}}$. Nonetheless, this method introduces a challenge: accurately computing ${{{\bf{\tilde{q}}}}_{1}}$, since it doesn’t inherently align with item popularity. To rectify this, we may simply mirror the calculation of ${\bf{q}_{1}}\leftarrow\frac{{\mathbf{Y}^{\top}}{{\bf{e}}}}{||{\mathbf{Y}^{\top}}{{\bf{e}}}||}$ to ${{{\bf{\tilde{q}}}}_{\bf{1}}}\leftarrow\frac{{\bf{V}}{{\bf{U}}^{\top}}{{\bf{e}}}}{||{\bf{V}}{{\bf{U}}^{\top}}{{\bf{e}}}||}$. This approach is clued by our Observation 1 and Theorem 1: a matrix’s principal singular vector tends to align with the column sum vector, especially when the vector showcases a long-tailed distribution. The discussions presented in Appendix B.3 validate the precision of this strategy. In essence, our ReSN optimizes the following loss function:
@@
4.2 Discussions
The proposed ReSN have the following merits: 1) Model-Agnostic. The proposed ReSN is model-agnostic and easy to implement. Given that ReSN introduces merely a regularization term, it can be easily plugged into existing embedding-based methods with minimal code augmentation. 2) Efficiency. The regularizer can be fast computed from right to left — it predominantly requires the multiplication of a $n\times d$ (or $m\times d$) matrix with a vector. With a time complexity of $O((n+m)d)$, ReSN is highly efficient. Appendix B.5 also provides empirical evidence. The additional time for calculating the regularizer is negligible. 3) Suitable for BPR Loss. As delineated in Section 2, while BPR can be regarded as a specialized point-wise loss, it involves the concept of hyper-items. Our theoretical analyses presented in Appendix A.1.2 demonstrate ReSN can be a logical regularizer even for the BPR loss.
@@ 4.2 Discussions
Compared with Other Regularizers: 1) recent studies [9, 69, 53] has employed regularizers to alleviate the dimensional collapse of user/item embeddings. Our ReSN diverges from these methods in two key aspects: Firstly, ReSN imposes constraints directly onto the prediction matrix, unlike the embedding matrix constraints utilized in these methods. This distinction is of significance due to the inherent spectral gap between the embeddings and the prediction matrix. Secondly, ReSN explicitly modulates the influence of the principal spectrum that captures popularity information, while these methods mainly focuses on promoting embedding uniformity. ReSN directly and solely mitigates the impact of the memorized popularity signal, thus demonstrating high efficacy in mitigating popularity bias; while others may disrupt the spectral structure of the prediction, potentially compromising model accuracy. 2) Other researchers have introduced various regularizers tailored to combat popularity bias [74, 36, 47]. However, these approaches are often heuristic, applying strong constraints to model predictions that may break the model’s original spectrum. While it could mitigate popularity bias, this approach may also impair the model’s ability to capture other useful signals, significantly compromising recommendation accuracy. Contrasting this, our ReSN is a light and theoretic-grounding approach — only modulates the influence of the principle spectrum.
5 Experiments
5.1 Experiment Settings
Datasets and Metrics. We adopt seven real-world datasets, Yelp2018 [27], Douban [50], Movielens [63], Gowalla [26], Globo [20], Yahoo!R3 [38] and Coat [49] for evaluating our model performance. Details about these datasets refer to C.1.
We adopt three representative testing paradigms for comprehensive evaluations: 1) Common: We employ the conventional testing paradigm in RS, wherein the datasets are randomly partitioned into training (70%), validation (10%), and testing (20%). We also report the accuracy-fairness trade-off in this setting. 2) Debiased: Closely referring to [56, 7, 71], we sample an debiased test set where items are uniformly distributed, aiming to evaluate the model’s efficacy in mitigating popularity bias. 3) Uniform-exposure: We also adopt the uniform exposure paradigm for model testing as the recent work [64]. Notably, the datasets Yahoo!R3 and Coat contain a small dataset collected through a random recommendation policy. Such data isolate the popularity bias from uneven exposure, offering a more precise estimation of user preferences. Consequently, we train our recommendation model on conventionally biased data and then test it on these uniformly-exposed data.
For evaluation metrics, we adopt the widely-used NDCG@K for evaluating accuracy [32]. We simply adopt $K=5$ for Yahoo and Coat datasets and $K=20$ for the other datasets as recent work [27, 64, 62]. We observe similar results with other metrics. We also employ the ratio of pop/unpopular items for illustrating the severity of popularity bias in recommendations. Here we closely refer to recent work [68] to define popular and unpopular items. We sort the items according to their popularity in descending order, and divide items into five groups ensuring the aggregated popularity of items within each group is the same. We define the items in the most popular groups as popular items, while the others as unpopular.
Baselines. The following methods are compared: 1) MACR (KDD’21 [56]), PDA (SIGIR’21 [68]): the representative causality-based debiasing methods, which posit a causal graph [42] for the recommendation procedure and leverage causal inference to mitigate popularity bias accordingly; 2) InvCF (WWW’23 [64]): the SOTA method that addresses popularity bias by disentangling the popularity from user preference. 3) Zerosum (Recsys’22 [47]) , IPL (SIGIR’23 [36]): the representative methods based on regularizers, which penalize the score differences or constrain the ratio of the predicted preference with the exposure.
@@ 5.1 Experiment Settings
Table 2: Performance comparison in terms of NDCG between ReSN and other baselines across seven datasets and three testing paradigms. The “Com”(refers to “Common”) represents the paradigm where the training and test datasets are partitioned randomly; “Deb”(refers to “Debiased”) represents the paradigm where a debiased test dataset is formulated based on item popularity; “Uni”(refers to “Uniform-exposure”) represents the paradigm where the test data is uniformly-exposed. The best result is bolded and the runner-up is underlined. The mark ‘*’ denotes the improvement achieved by ReSN over best baseline is significant with $p<0.05$.
@@
Figure 3: Pareto curves of compared methods illustrating the trade-off between accuracy and fairness.
Figure 4: NDCG@20 comparison with methods for addressing Dimension Collapse.

### v6

@@ 3.2. Popularity Bias Amplification Effect
As the model embedding dimension $d$ is reduced, the relative prominence of the principal singular value increases ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}\sigma_{k}^{2}\uparrow$), and the recommendation increasingly favors popular items.
@@
3.2.2 Theoretical Analyses. In this subsection, we focus on establishing a theoretical relationship between singular values and the ratio of popular items in recommendations. For readers interested in the theoretical support of the impact of gradient optimization, we refer them to the Appendix A.3, which are relatively straightforward by invoking recent gradient theory (Chou et al., 2023; Saxe et al., 2019). For convenience, our analysis here concentrates on the ratio of the most popular item in top-1 recommendations. We have:
@@
(3) $\small\eta\geq\frac{1}{n}\phi\left({\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}(\frac{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}{{{\sigma_{1}}}}-1)}\right)$
@@
where $\phi(a)=\sum_{u\in\mathcal{U}}{\mathbf{I}[p_{1u}>a]}$ is an inverse cumulative function calculating the number of elements $p_{1u}$ in the left principal singular vector $\mathbf{p}_{1}$ exceeding a given value $a$, and the function $\mathbf{I}[.]$ signifies an indicator function.
@@
The detailed proof is available in Appendix A.2. This theorem vividly showcases the influence of dimension reduction on popularity bias. Essentially, as dimension reduction intensifies the relative prominence of the principle singular value ($\frac{{{\sigma_{1}}}}{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}\uparrow$), the input of the function $\phi(.)$ decreases ($\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}(\frac{{\sum\nolimits_{1\leq k\leq L}{{\sigma_{k}}}}}{{{\sigma_{1}}}}-1)\downarrow$). Given the monotonically decreasing nature of $\phi(.)$, dimension reduction thus escalates the ratio of most popular items in recommendations. Interestingly, the theorem illustrates the impact of a long-tailed distribution on popularity bias. A larger $\alpha$ (indicating a more skewed item popularity distribution) decreases the value of $\frac{{\sqrt{2\zeta(2\alpha)}}}{{1-{2^{-\alpha}}}}$, further elevating the lower bound of the ratio, intensifying bias.
4. Proposed Method
In this section, we first introduce our proposed debiasing method, followed by a discussion of its properties and a comparison with other debiasing approaches.
4.1. ReSN: Regulation with Spectral Norm
@@
The above analyses elucidate the essence of the popularity bias amplification — the undue influence of the principal spectrum. Therefore, the core of an effective debiasing strategy naturally lies on mitigating this effect. To address this, we propose ReSN which leverages Spectral Norm Regularizer to penalize the magnitude of principal singular value:
(4) ${\mathcal{L}_{{\mathop{\rm Re}\nolimits}SN}}={\mathcal{L}_{R}}({\mathbf{Y}},{\mathbf{\hat{Y}}})+\beta||{\mathbf{\hat{Y}}}||_{2}^{2}\vspace{-2pt}$
@@
where ${\mathcal{L}_{R}}({\mathbf{Y}},{\mathbf{\hat{Y}}})$ is original recommendation loss, and ||.$||.||_{2}$ denote the spectral norm of a matrix measuring its principle singular value; $\beta$ controls the contribution from the regularizer.
@@
However, there are practical challenges: 1) the $n\times m$ dimensional matrix $\mathbf{\hat{Y}}$ can become exceptionally large, often comprising billions of entries, making direct calculations computationally untenable; 2) Existing methods to determine the gradient of the spectral norm are iterative (Yoshida and Miyato, 2017; Sun et al., 2021), which further adds computational overhead.
To circumvent these challenges, we make two refinements:
Firstly, given the alignment of the principal singular vector $\mathbf{q}_{1}$ with item popularity $\mathbf{r}$, the calculating of the spectral norm can be simplified into: $||{\mathbf{\hat{Y}}}||_{2}^{2}=||{\mathbf{\hat{Y}}}{{\mathbf{q}}_{1}}||^{2}\approx||{\mathbf{\hat{Y}r}}||^{2}/||\mathbf{r}||^{2}$, where ||.||||.||| | . | | denotes the L2-norm of a vector. It transforms the calculation of the complex spectral norm of a matrix to a simple L2-norm of a vector, avoiding iterative algorithms by leveraging the singular vector property. Further, the item popularity $\mathbf{r}$ can be quickly computed via ${\mathbf{r}}={\mathbf{Y}^{\top}}{{\mathbf{e}}}$, where ${\mathbf{e}}$ represents a $n$-dimension vector filled with ones.
@@
Secondly, we exploit the low-rank nature of the matrix $\mathbf{\hat{Y}}$. For models based on embeddings, $\mathbf{\hat{Y}}$ can be expressed as ${\mathbf{\hat{Y}}}=\mu({\mathbf{U}}{{\mathbf{V}}^{\top}})$, where ${\mathbf{U}}$ and ${{\mathbf{V}}}$ represent the embeddings associated with users and items, respectively, and $\mu(.)$ designates an activation function. Our approach turns to penalize the spectral norm of the matrix before the introduction of the activation function. This is motivated by the ease of computation: $||{\mathbf{U}}{{\mathbf{V}}^{\top}}||_{2}^{2}=||{\mathbf{U}}({{\mathbf{V}}^{\top}}{{{\mathbf{\tilde{q}}}}_{1}})||^{2}$, where ${{{\mathbf{\tilde{q}}}}_{1}}$ denotes the right principal vector of the matrix ${\mathbf{U}}{{\mathbf{V}}^{\top}}$. By adopting this method, we circumvent the computationally-intensive task of processing the entire matrix $\mathbf{\hat{Y}}$. Nonetheless, this method introduces a challenge: accurately computing ${{{\mathbf{\tilde{q}}}}_{1}}$, since it doesn’t inherently align with item popularity. To rectify this, we may simply mirror the calculation of ${\mathbf{q}_{1}}\leftarrow\frac{{\mathbf{Y}^{\top}}{{\mathbf{e}}}}{||{\mathbf{Y}^{\top}}{{\mathbf{e}}}||}$ to ${{{\mathbf{\tilde{q}}}}_{\mathbf{1}}}\leftarrow\frac{{\mathbf{V}}{{\mathbf{U}}^{\top}}{{\mathbf{e}}}}{||{\mathbf{V}}{{\mathbf{U}}^{\top}}{{\mathbf{e}}}||}$. This approach is clued by our Observation 1 and Theorem 1: a matrix’s principal singular vector tends to align with the column sum vector, especially when the vector showcases a long-tailed distribution.
To empirically validate the accuracy and rationality of the proposed method, we computed the ideal value of $||{\mathbf{U}\mathbf{V}^{\top}}||_{2}^{2}$, as well as the estimated $\frac{||{\mathbf{U}}{{\mathbf{V}}^{\top}}{\mathbf{V}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}{{||{\mathbf{U}\mathbf{V}^{\top}{\mathbf{e}}}||^{2}}}$ from ReSN, training the MF model with two losses on three datasets. The results are shown in the Table 2. According to the table, we found that the actual spectral norms and our approximate estimates are very close across diverse losses and datasets.This indicates that the singular vector $\mathbf{\tilde{q}}_{1}$ obtained through $\frac{\mathbf{U}\mathbf{V}^{\top}\mathbf{e}}{||\mathbf{U}\mathbf{V}^{\top}\mathbf{e}||}$, serves as an accurate surrogate for the true value of $\mathbf{q}_{1}$. Therefore, the estimated regularization term is a accurate surrogate for the spectral norm $||{\mathbf{U}\mathbf{V}^{\top}}||^{2}_{2}$ which validates the precision of this strategy.
In essence, our ReSN optimizes the following loss function:
@@
4.2. Discussions
The proposed ReSN have the following aspects:
Model-Agnostic: The proposed ReSN is model-agnostic and easy to implement. Given that ReSN introduces merely a regularization term, it can be easily plugged into existing embedding-based methods with minimal code augmentation.
Efficiency: The regularizer can be fast computed from right to left — it predominantly requires the multiplication of a $n\times d$ (or $m\times d$) matrix with a vector. With a time complexity of $O((n+m)d)$, ReSN is highly efficient. Section 5.5 also provides empirical evidence. The additional time for calculating the regularizer is negligible.
Suitable for BPR Loss: As delineated in Section 2, while BPR can be regarded as a specialized point-wise loss, it involves the concept of hyper-items. It means that the regularizer should be conducted on the embedding matrix of hyper-items $V^{\prime}\in\mathbb{R}^{m^{2}\times d}$, i.e., $\frac{||{\mathbf{U}}{{\mathbf{V^{\prime}}}^{\top}}{\mathbf{V^{\prime}}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}{||{\mathbf{V^{\prime}}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}$, rather than $\frac{||{\mathbf{U}}{{\mathbf{V}}^{\top}}{\mathbf{V}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}{||{\mathbf{V}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}$. In the following, we will build their approximations. For the numerator part, we have:
${{{\mathbf{V^{\prime}}}}^{\top}}{\mathbf{V^{\prime}}}$ $=\sum\limits_{i,j\in I}{({{\mathbf{v}}_{i}}-{{\mathbf{v}}_{j}})^{\top}{{({{\mathbf{v}}_{i}}-{{\mathbf{v}}_{j}})}}}=2m\mathbf{V}^{\top}{\mathbf{V}}-2m^{2}{\mathbf{\bar{v}}}^{\top}{{{\mathbf{\bar{v}}}}}$
where $\bar{\mathbf{v}}=\sum_{i\in I}{{{\mathbf{v}}_{i}}}/m$ denote the mean vector of the item embeddings. Furthermore, current literature posits that an ideal item representation should emulate a uniform distribution over the unit ball (Wang et al., 2022). This implies that $\bar{\mathbf{v}}$ tends to gravitate towards the origin. Thus, ${{{\mathbf{V^{\prime}}}}^{\top}}{\mathbf{V^{\prime}}}$ can be approximated by $\mathbf{V}^{\top}{\mathbf{V}}$ and $||{\mathbf{U}}{{\mathbf{V}^{\prime}}^{\top}}{\mathbf{V}^{\prime}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}$ can be approximated by $||{\mathbf{U}}{{\mathbf{V}}^{\top}}{\mathbf{V}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}$.
Similarly, for the denominator:
$||{\mathbf{V^{\prime}U^{\top}e}}||^{2}$ $=\sum\limits_{i,j\in I}{({{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}}-{{\mathbf{v}}_{j}}{\mathbf{U^{\top}e}})^{\top}{{({{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}}-{{\mathbf{v}}_{j}}{\mathbf{U^{\top}e}})}}}$
$=2m\sum\limits_{i\in I}({{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}})^{\top}({{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}})-2(\sum\limits_{i\in I}{{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}})^{\top}(\sum\limits_{i\in I}{{\mathbf{v}}_{i}}{\mathbf{U^{\top}e}})$
$=2m({\mathbf{VU^{\top}e}})^{\top}({\mathbf{VU^{\top}e}})-2m^{2}({{\mathbf{\bar{v}}}}{\mathbf{U^{\top}e}})^{\top}({{\mathbf{\bar{v}}}}{\mathbf{U^{\top}e}})$
We can deduce ${||{\mathbf{V}^{\prime}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}$ can be approximated by ${||{\mathbf{V}}{{\mathbf{U}}^{\top}}{\mathbf{e}}||^{2}}$. Consequently, ReSN emerges as a logical regularizer even for the BPR loss. This assertion is also validated by our experiments.
Table 2. Comparison between the actual spectral norm and the estimated approximation.
Datasets MSE BCE
$||{\mathbf{UV^{\top}}}||_{2}^{2}$ $||{\mathbf{U(V^{\top}{{\mathbf{\tilde{q}}}_{1}})}}||^{2}$ $||{\mathbf{UV^{\top}}}||_{2}^{2}$ $||{\mathbf{U(V^{\top}{{\mathbf{\tilde{q}}}_{1}})}}||^{2}$
@@ 4.2. Discussions
Differences from Methods on Dimensional Collapse: Recent studies (Chen et al., 2023b; Zhang et al., 2023b; Wang et al., 2022) has also employed regularizers to alleviate the dimensional collapse of user/item embeddings. Our ReSN diverges from these methods in two key aspects: 1) ReSN imposes constraints directly onto the prediction matrix, unlike the embedding matrix constraints utilized in these methods. This distinction is of significance due to the inherent spectral gap between the embeddings and the prediction matrix. 2) ReSN explicitly modulates the influence of the principal spectrum that captures popularity information, while these methods mainly focuses on promoting embedding uniformity. ReSN directly and solely mitigates the impact of the memorized popularity signal, thus demonstrating high efficacy in mitigating popularity bias; while others may disrupt the spectral structure of the prediction, potentially compromising model accuracy.
Differences from Regularization-based Debiasing methods: Various regularizers are introduced to combat popularity bias (Zhu et al., 2021; Liu et al., 2023; Rhee et al., 2022; Zhang et al., 2023b). However, except (Zhang et al., 2023b) as discussed before, existing approaches are typically heuristic, applying strong constraints to model predictions that may break the model’s original spectrum. While it could mitigate popularity bias, this approach may also impair the model’s ability to capture other useful signals, significantly compromising recommendation accuracy. Contrasting this, our ReSN is a light and theoretic-grounding approach — it motivated by the core reason of bias amplification and only modulates the influence of the principle spectrum.
@@
Table 3. Performance comparison in terms of NDCG between ReSN and other baselines across seven datasets and three testing paradigms. The “Com”(refers to “Common”) represents the paradigm where the training and test datasets are partitioned randomly; “Deb”(refers to “Debiased”) represents the paradigm where a debiased test dataset is formulated based on item popularity; “Uni”(refers to “Uniform-exposure”) represents the paradigm where the test data is uniformly-exposed. The best result is bolded and the runner-up is underlined. The mark ‘*’ denotes the improvement achieved by ReSN over best baseline is significant with $p<0.05$.
@@
5. Experiments
We conduct experiments to address the following questions:
RQ1: How does ReSN perform compared with other methods?
RQ2: Is ReSN suitable for diversified loss functions and backbones?
RQ3: What is the impact of regularizer coefficient $\beta$?
RQ4: How is the efficiency of ReSN ?
5.1. Experiment Settings
Datasets and Metrics. We adopt seven real-world datasets, Yelp2018 (He et al., 2020), Douban (Song et al., 2019), Movielens (Yu and Qin, 2020), Gowalla (He et al., 2017), Globo (de Souza Pereira Moreira et al., 2018), Yahoo!R3 (Marlin and Zemel, 2009) and Coat (Schnabel et al., 2016) for evaluating our model performance. Details about these datasets refer to Appendix C.1.
We adopt three representative testing paradigms for comprehensive evaluations: 1) Common: We employ the conventional testing paradigm in RS, wherein the datasets are randomly partitioned into training (70%), validation (10%), and testing (20%). We also report the accuracy-fairness trade-off in this setting. 2) Debiased: Closely referring to (Wei et al., 2021; Bonner and Vasile, 2018; Zheng et al., 2021), we sample an debiased test set where items are uniformly distributed, aiming to evaluate the model’s efficacy in mitigating popularity bias. 3) Uniform-exposure: We also adopt the uniform exposure paradigm for model testing as the recent work (Zhang et al., 2023a; Wang et al., 2024a; Lin et al., 2024). Notably, the datasets Yahoo!R3 and Coat contain a small dataset collected through a random recommendation policy. Such data isolate the popularity bias from uneven exposure, offering a more precise estimation of user preferences. Consequently, we train our recommendation model on conventionally biased data and then test it on these uniformly-exposed data.
For evaluation metrics, we adopt the widely-used NDCG@K for evaluating accuracy (Krichene and Rendle, 2020). We simply adopt $K=5$ for Yahoo and Coat datasets and $K=20$ for the other datasets as recent work (He et al., 2020; Zhang et al., 2023a; Yu et al., 2023). We observe similar results with other metrics. We also employ the ratio of pop/unpopular items for illustrating the severity of popularity bias in recommendations. Here we closely refer to recent work (Zhang et al., 2021) to define popular and unpopular items. We sort the items according to their popularity in descending order, and divide items into five groups ensuring the aggregated popularity of items within each group is the same. We define the items in the most popular groups as popular items, while the others as unpopular.
Baselines. The following methods are compared: 1) MACR (KDD’21 (Wei et al., 2021)), PDA (SIGIR’21 (Zhang et al., 2021)): the representative causality-based debiasing methods, which posit a causal graph (Pearl, 2009) for the recommendation procedure and leverage causal inference to mitigate popularity bias accordingly; 2) InvCF (WWW’23 (Zhang et al., 2023a)): the SOTA method that addresses popularity bias by disentangling the popularity from user preference. 3) Zerosum (Recsys’22 (Rhee et al., 2022)) , IPL (SIGIR’23 (Liu et al., 2023)): the representative methods based on regularizers, which penalize the score differences or constrain the ratio of the predicted preference with the exposure.
Table 4. NDCG@20 comparison with methods for addressing Dimension Collapse under the debiased testing paradigm.